OpenStudy (automaticloveletter):

A system of equations is shown below: y = 8x − 2 y = 9x − 7 What is the solution to the system of equations? (−5, 38) (−5, −38) (5, 38) (5, −38)

pooja195 (pooja195):

Substitution would be the easiest here since its given what y equals 8x-2=9x-7 I plugged in equation 1 into equation 2 to find the x value :) 8x-2=9x-7 Please solve for x

OpenStudy (automaticloveletter):

5

pooja195 (pooja195):

yes now we plug that into equation 1 y = 8(5)− 2 solve for y

OpenStudy (automaticloveletter):

38?

pooja195 (pooja195):

yes :) so which answer choices mathces?

pooja195 (pooja195):

*matches

OpenStudy (automaticloveletter):

C

pooja195 (pooja195):

good